OpenStudy (anonymous):

Let u = <-5, -3>, v = <-6, -1>. Find -3u + 5v. @SolomonZelman

OpenStudy (solomonzelman):

I hope this is not about vectors :)

OpenStudy (anonymous):

i only have like 3 more problems left that i need help in

OpenStudy (anonymous):

oh noo is vectors bad?

OpenStudy (solomonzelman):

u ≤ -5 v ≤ -6 this is the given?

OpenStudy (solomonzelman):

Yes, vectors are very bad. I am a trig student, and the only think I didn't understand from calc1 book was this section about vectors.

OpenStudy (solomonzelman):

only *thing

OpenStudy (anonymous):

hahaha thats okay, but thanks for the help with everything else (:

OpenStudy (anonymous):

\[ -3\mathbf u = -3\langle -5,-3\rangle = \langle -5\cdot -3,-3\cdot -3\rangle = \langle 15,9\rangle \]

OpenStudy (anonymous):

Do the same to find \(5\mathbf v\).

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

5(-6,-1)= (-30,-5)

OpenStudy (anonymous):

i am right> @wio ??

OpenStudy (anonymous):

<-15, 4>

OpenStudy (johnweldon1993):

That is the correct <v> yes so now <15 , 9> + <-30 , -5> = <15 - 30, 9 - 5> = Yes <-15 , 4>